OpenStudy (anonymous):

Select next number in series: 298 209 129 58 4 (My answer is a. -13 b. -93 c. -66 d. -57

OpenStudy (anonymous):

Hi AS, pls check again and inform if the last number in the series is 4 or -4 ???

OpenStudy (anonymous):

last number is 4. thank you

OpenStudy (anonymous):

the differences are 298 (89) 209 (80) 129 (71) 58 (54) 4, but i can't find a pattern there.

OpenStudy (anonymous):

That is why I asked if the last number was -4 or not

OpenStudy (anonymous):

My original answer was -13. But then my friend came up with -66. Think -66 is correct as the spread is 9

OpenStudy (anonymous):

If it is -4 then it works out Each time the number being subtracted decreases by 9

OpenStudy (anonymous):

So 58 - 62 = -4 and -4 - 53 = -57

OpenStudy (anonymous):

That looks right on everything, but for the fact that the last number of the actual sequence doesn't fit.

OpenStudy (anonymous):

AS, how did yr friend land up with -66?? what's the working??

OpenStudy (anonymous):

For all u know, it might be a typo!!

OpenStudy (anonymous):

exactly, twice though? i doubt that.

OpenStudy (anonymous):

Not from AS, it might be a typo in the printed paper he got

OpenStudy (anonymous):

Harkirat. Went back to the math problem. You are very sharp. The last number is a -4. So sorry. Does it make better sense now.

OpenStudy (anonymous):

Sure does :)

OpenStudy (anonymous):

Do I get two medals for that????

OpenStudy (anonymous):

Yes, I gave you the medals...very well deserved. So what is your final answer? Like i said, originally I had -13 but got lost somewhere

OpenStudy (anonymous):

Final answer is -57

OpenStudy (anonymous):

Harkirat: Please explain -57. the differences are 298 (89) 209 (80) 129 (71) 58 (54) -4,

OpenStudy (anonymous):

theres a change, because the final is -4, it is 62, not 54, so the difference decreases by 9 between each set, therefore the next one is 62-9=53, so -4-53=-57

OpenStudy (anonymous):

See, it goes like this 298 - 89 = 209 209 - 80 = 129 129 - 71 = 58 58 - 62 = -4 -4 -53 = -57 Got it now??

OpenStudy (anonymous):

Thank you so much. Yes, I see it now. You both Harkirat & buksbelly are very, very, sharp. Enjoy the rest of your evening. It's after 1:00 am., in California. Goodnight

OpenStudy (anonymous):

Bonne nuit!

OpenStudy (anonymous):

Shubh Ratri (Good Night in Hindi )